OpenStudy (anonymous):

Will give a medal to anyone that answers plz 6. simplify x / 7x+x^2 7. simplify -14x^3 / x^3-5x^4 8. x-2 / x^2+4x-12 9.x^2+3x-4 / x+4 10. 2/3a by 2/a^2 11. x+2/5x-15 by (10x^2-25x-15) 12. (x^2-9 / x-5)/x+3

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

and i will fan you if u answer it

OpenStudy (anonymous):

6. \(\Large \frac{x}{7x+x^2}=\frac{x}{x(7+x)}=\frac{1}{7+x}\)

OpenStudy (anonymous):

7.\(\Large \frac{-14x^3}{x^3-5x^4} =\frac{-14x^3}{x^3(1-5x)}=\frac{-14}{1-5x} \)

OpenStudy (anonymous):

8. \(\Large \frac{x-2}{x^2+4x-2}=\frac{x-2}{(x-2)(x+6)}=\frac{1}{x+6}\)

OpenStudy (anonymous):

9. \(\Large \frac{x^2+3x-4}{x+4}=\frac{(x+4)(x-1)}{x+4}=x-1\)

OpenStudy (anonymous):

10. \(\huge \frac{\frac{2}{3a}}{\frac{2}{a^2}}=\frac{a^2}{3a}=\frac{a}{3}\)

OpenStudy (anonymous):

11. \(\huge \frac{\frac{x+2}{5x-15}}{10x^2-25x-15}=\frac{x+2}{(5x-15)(10x^2-25x-15)}\\= \huge \frac{x+2}{(5(x-3))((x-3)(2x+1))}=\frac{x+2}{5(x-3)^2(2x+1)}\)

OpenStudy (anonymous):

12. \(\huge \frac{\frac{x^2-9}{x-5}}{x+3} =\frac{x^2-9}{(x-5)(x+3)}\\ \huge =\frac{(x-3)(x+3)}{(x-5)(x+3)}=\frac{x-3}{x-5}\)